# Regularization methods for solving hierarchical variational inequalities with complexity guarantees

### v1

@@ 3.3 Combined Complexity Analysis
We assume $\mathsf{Z}_{n}$ is domain-forward on $\operatorname{dom}(\mathsf{M})$, namely that $\mathsf{Z}_{n}(\operatorname{dom}(\mathsf{M}))\subset\operatorname{dom}(\mathsf{M})$. Finally, we assume that $\theta_{n,k}$ is uniformly lower bounded by a positive quantity, such that $\overline{Q}_{n}\leq\overline{Q}<1$.
@@
Assume the setting of Corollary 3.3 and that Assumptions 2, 3, 4 and 6 hold. Algorithm 2 (DANTE) reaches a point $\overline{w}_{N}$ satisfying the inequalities in Theorem 3.8 after at most
$K=\frac{N(2\log(C)+\log(\overline{Q}))+2\sum_{n=1}^{N}\log(1/\varepsilon_{n})}{\log(1/\overline{Q})}$
inner iterations. If $\varepsilon_{n}=(n+1)^{-\eta}$ where $\eta\geq b$, we reach the desired point $\overline{w}_{N}$ after at most ${\mathcal{O}}(N\log N)$ inner iterations.
@@
By Corollary 3.3, each outer loop $n\in\{0,\ldots,N-1\}$ runs at most
$\overline{K}_{n}=\left\lceil\frac{2\log(C)+2\log(1/\varepsilon_{n})}{\log(1/\overline{Q})}\right\rceil\leq\frac{2\log(C)+2\log(1/\varepsilon_{n})}{\log(1/\overline{Q})}+1$
inner iterations. Specifically, the inequalities of Theorem 3.8 are satisfied after at most
$K=\sum_{n=0}^{N-1}\overline{K}_{n}\leq\frac{2N\log(C)+2\sum_{n=1}^{N}\log(1/\varepsilon_{n})}{\log(1/\overline{Q})}+N=\mathcal{O}\left(N+\sum_{n=0}^{N-1}\log(1/\varepsilon_{n})\right)$
By setting $\varepsilon_{n}=(n+1)^{-\eta}$, we obtain that $\sum_{n=0}^{N-1}\log(1/\varepsilon_{n})\leq\eta N\log(N)$, as wanted. $\blacksquare$
Remark 3.5.
Theorem 3.11 shows that accounting for the total number of inner iterations, rather than only the number of outer iterations, introduces merely a logarithmic factor. This is expected, as double-loop schemes inherently incur such a term. Consequently, the complexities stated in Corollaries 3.9 and 3.10 are preserved, up to logarithmic factors, when measured in terms of total inner iterations.
@@ 4 Relevant Splitting Methods
With the following examples it will become clear that the fixed-point encoding $\mathsf{T}^{(w,\beta)}_{k}$ not only depends on the problem parameters $\alpha$ and $\beta$ but also on the step size $\gamma_{k}$. Therefore, we treat the step size policy is treated in our approach as an ingredient of the fixed-point encoding strategy and consequently incorporated in the map $\mathsf{T}^{(w,\beta)}_{k}$, rather than being a parameter of the inner loop presented in Algorithm 1.
@@ 4.1 Two-Operator Splitting Schemes
    The lower level solution set $\mathcal{S}_{0}$ is nonempty;
@@
We note that $\operatorname{Fix}(\mathsf{T}^{(w,\beta)}_{\gamma})=\operatorname{Zer}(A+\Phi_{\alpha,\beta}^{w})$, so $\mathsf{Z}=\operatorname{Id}$ is the fixed-point transportation map. This is the setting studied in [marschner_tikhonov_2025].
@@
Pick $z_{1},z_{2}\in\mathcal{H}$. By nonexpansiveness of $\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}$, and $L_{\alpha,\beta}$-Lipschitz continuity and $\alpha$-strong monotonicity of $\Phi_{\alpha,\beta}$,
@@
In a similar way, we can employ the Backward-Forward splitting [attouch_backward_2018] for the auxiliary problem. Consider the operator
@@
Note that $\operatorname{Zer}(\mathsf{A}+\Phi_{\alpha,\beta}^{w})=\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}(\text{Fix}(\mathsf{T}^{(w,\beta)}_{\gamma}))$, which means that the last iterate of the inner loop scheme needs one more evaluation of the resolvent of the operator $\mathsf{A}$ in order to get close to the unique zero of $\mathsf{A}+\Phi_{\alpha,\beta}^{w}$. Specifically, the fixed-point transportation map is $\mathsf{Z}=\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}$.
@@
Assume Assumption 7 holds and let $\mathsf{T}_{\gamma}$ be the fixed-point encoding defined in (BF). It holds that $\mathsf{T}_{\gamma}$ is $q$-Lipschitz where
@@ 5 Numerical Experiments
All the code may be found on the author’s GitHub page [^1] See https://github.com/Hierarchical-VIs/Regularisation-Methods-for-HVIs., and is run on Intel Xeon Platinum 8380 CPUs.
@@ 5.1 Equilibrium selection
This matches Problem (P) with $\mathsf{G}=\nabla\phi$ for the upper level operator, and $\mathsf{M}=\mathsf{F}+\operatorname{\mathsf{NC}}_{X}$ for the lower level operator.
@@ Results.
Moreover, we note that the optimality and feasibility gaps may be computed through a small subproblem in this simple example. Figures 3 and 4 plot these values, for the different initial points. Although we observe differences, we note that all curves are decreasing to 0. We note the importance of the tuning of the proximal parameter $\alpha$. In the present example, a smaller value yields faster convergence. As will be evidenced in later simulations, this is not always the case.
@@ 5.2 Least-Norm Least-Squares
We test the algorithm on four randomly generated instances with different dimensions given by $(P,Q)\in\{(70,100),(100,200),(100,500),(300,500)\}$. For each instance, the matrix $A$ is generated as a low-rank (and sparse) matrix via $A=U_{1}\cdot U_{2}$, where $U_{1}\in\mathbb{R}^{P\times R}$ and $U_{2}\in\mathbb{R}^{R\times Q}$ are chosen such that each component follows a standard normal law, with $R=50$ to induce a low-rank structure. Moreover, to ensure the eigenvalues are not too large such that the step-size is not too small, we clip each singular value to the interval $[0,10]$. A vector $s\in\mathbb{R}^{Q}$ is randomly generated with $20$ non-zero entries according to a uniform law on $[0,10]$, and the observation vector $b$ is constructed as $b=As+\nu$, where $\nu$ is a normal random vector with small entries, namely a standard normal vector scaled by a factor $0.1$. During the random generation we ensure that $z=A^{\dagger}b\in X$, such that the analytical solution is known to be $z$, where $A^{\dagger}$ is the Moore-Penrose inverse.
@@ Results.
Figures 7 and 8 show the evolution of the error sequence $\|w_{n}-z\|$ and the lower-level function gap sequence $\tfrac{1}{2}\|Aw_{n}-b\|^{2}-\tfrac{1}{2}\|Az-b\|^{2}$ across iterations. As expected, we observe a decrease for each problem. We note that the initial decrease for the error sequence is consequential and stagnates for a number of iterations afterwards. In fact, for $\alpha=0.1$, the prescribed $2000$ iterations are not enough to observe a second decrease phase, whereas it is sufficient for $\alpha\in\{1,10\}$.
@@ 5.3 Image Inpainting
For the inner loop, we set $\theta_{k}\equiv\theta=0.75$ and we consider $\tau_{k}\equiv\tau$ to be constantly the largest value satisfying Equation (3.1) across every outer iteration. We set $\beta_{n}=(n+1)^{-\eta}$ for $\eta=0.55$ and $\varepsilon_{n}=\overline{\varepsilon}\cdot(n+1)^{-2}$ for $\overline{\varepsilon}=2$. We run a total of $5000$ iterations, with an initial point $w_{0}=Y_{\text{corrupt}}$. We set the regularization parameter to be $\sigma=50$, and generate $\Omega$ randomly such that $20\%$ of the image is corrupted.
@@ Results.
Figure 9 shows the original image, the corrupt image and the image restored using the above procedure for various proximal parameters. Table 1 reports the total number of inner iterations run, along with the time spent to produce the results. Figure 10 plots the lower-level objective across outer iterations for each value of $\alpha$, which is expected to converge to the lower-level objective value and not to 0, as seen in the plot. We observe a loss in the solution quality for $\alpha=100$, although it is the fastest method. We also note that the solution qualities are very similar for $\alpha=0.1$ and $\alpha=1$, but the latter requires a larger number of total iterations.
@@
Figure 10: Evolution of lower-level objective function $\tfrac{1}{2}\|\mathcal{R}(Y)-Y_{\text{corrupt}}\|^{2}+\sigma\|Y\|_{*}$ for various proximal parameters.
@@ 6 Conclusion
In this work, we proposed a double loop path following method for solving a hierarchical variational inequalities. The inner loop employs a Krasnoselskii-Mann type iteration for solving a suitable auxiliary problem. The outer loop takes the form of a restarting procedure, giving rise to a diagonal equilibrium tracking method (DANTE). The obtained convergence rates are comparable with the state of the art results, obtained for specific numerical schemes. Our analysis is very flexible with respect to the chosen fixed-point encoding strategy of the lower level solution. Importantly, our rates are independent of the specific fixed-point encoding, as long as the stated assumptions are satisfied. We belief that our work opens up the possibilities for various interesting directions for future research. First, it would be very interesting to device a fully adaptive scheme without restarts. Second, an important future extension is the inclusion of stochastic data. All these directions are the subject of ongoing research.
Acknowledgments. Daniel Cortild acknowledges the support of the Clarendon Funds Scholarships. Meggie Marschner and Mathias Staudigl acknowledge financial support from the PGMO and the Deutsche Forschungsgemeinschaft (DFG) - Projektnummer 556222748 "non-stationary hierarchical minimization".

### v2

@@ 3.3 Total Complexity Analysis
We assume $\mathsf{Z}_{n}$ is domain-forward on $\operatorname{dom}(\mathsf{M})$:
$\mathsf{Z}_{n}(\operatorname{dom}(\mathsf{M}))\subset\operatorname{dom}(\mathsf{M}).$
Finally, we assume that $\theta_{n,k}$ is uniformly lower bounded by a positive quantity, such that $\overline{Q}_{n}\leq\overline{Q}<1$.
@@
Let Assumptions 1-6 hold. Given $N\geq 1$, and inner loop tolerance sequence $(\varepsilon_{n})_{n=0}^{N}$, we have
$\mathcal{C}(N,(\varepsilon_{n})_{n=0}^{N})\leq\frac{N(2\log(C)+\log(\overline{Q}))+2\sum_{n=1}^{N}\log(1/\varepsilon_{n})}{\log(1/\overline{Q})},$
where the constant $C$ is defined in (3.2) and $\overline{Q}$ is defined in Assumption 3.
Specifically, if $\varepsilon_{n}=(n+1)^{-\eta}$ and $\beta_{n}=(n+1)^{-b}$ where $\eta\geq b$, then $\mathcal{C}(N,(\varepsilon_{n})_{n=0}^{N})={\mathcal{O}}(N\log N)$.
@@
By Corollary 3.3, for each restarting epoch $n\in\{0,\ldots,N-1\}$ the total number of iterations of the inner loop is bounded by
$\mathbb{K}(\varepsilon_{n})\leq\overline{K}_{n}=\left\lceil\frac{2\log(C)+2\log(1/\varepsilon_{n})}{\log(1/\overline{Q})}\right\rceil\leq\frac{2\log(C)+2\log(1/\varepsilon_{n})}{\log(1/\overline{Q})}+1.$
Hence,
$\sum_{n=0}^{N-1}\overline{K}_{n}\leq\frac{2N\log(C)+2\sum_{n=1}^{N}\log(1/\varepsilon_{n})}{\log(1/\overline{Q})}+N=\mathcal{O}\left(N+\sum_{n=0}^{N-1}\log(1/\varepsilon_{n})\right)$
gives an upper bound on the total number of iterations for Algorithm 2 for $N\geq 1$ restarts. By setting $\varepsilon_{n}=(n+1)^{-\eta}$, we obtain that $\sum_{n=0}^{N-1}\log(1/\varepsilon_{n})\leq\eta N\log(N)$. $\blacksquare$
Theorem 3.13 shows that accounting for the total number of inner iterations, rather than only the number of outer iterations, introduces merely a logarithmic factor. This is expected, as double-loop schemes inherently incur such a term. Consequently, the complexities stated in Corollaries 3.11 and 3.12 for the gap functions remain intact, up to logarithmic factors, when accounting for the total number of inner iterations.
@@ 4 Relevant Splitting Methods
With the following examples, it will become clear that the fixed-point encoding $\mathsf{T}^{(w,\beta)}_{k}$ not only depends on the problem parameters $\alpha$ and $\beta$ but also on step size sequences $(\gamma_{k})_{k}$. Therefore, we treat the step size policy as an ingredient of the fixed-point encoding strategy and consequently incorporate it in the map $\mathsf{T}^{(w,\beta)}_{k}$, rather than treating it as a parameter of the inner loop presented in Algorithm 1.
@@ 4.1 Two-Operator Splitting Schemes
    The lower-level solution set $\mathcal{S}_{0}$ is nonempty;
@@
We note that $\operatorname{Fix}(\mathsf{T}^{(w,\beta)}_{\gamma})=\operatorname{Zer}(\mathsf{A}+\Phi_{\alpha,\beta}^{w})$, so $\mathsf{Z}=\operatorname{Id}$ is the fixed-point transportation map. This is the setting studied in [marschner_tikhonov_2025].
@@
Pick $z_{1},z_{2}\in\mathcal{H}$ and set $T=T^{(w,\beta)}_{\gamma}$. By nonexpansiveness of $\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}$, and $L_{\alpha,\beta}$-Lipschitz continuity and $\alpha$-strong monotonicity of $\Phi_{\alpha,\beta}$,
@@
Similarly, we can employ the backward-forward splitting [attouch_backward_2018] for the auxiliary problem. Consider the operator
@@
Note that $\operatorname{Zer}(\mathsf{A}+\Phi_{\alpha,\beta}^{w})=\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}(\text{Fix}(\mathsf{T}^{(w,\beta)}_{\gamma}))$, which means that the last iterate of the inner loop scheme needs one more evaluation of the resolvent of the operator $\mathsf{A}$ to get close to the unique zero of $\mathsf{A}+\Phi_{\alpha,\beta}^{w}$. Specifically, the fixed-point transportation map is $\mathsf{Z}=\operatorname{\mathsf{J}}_{\gamma\mathsf{A}}$.
@@
Assume Assumption 7 holds, and let $\mathsf{T}_{\gamma}$ be the fixed-point encoding defined in (BF). It holds that $\mathsf{T}_{\gamma}$ is $q$-Lipschitz where
@@ 5 Numerical Experiments
All the code may be found on the author’s GitHub page [^1] See https://github.com/Hierarchical-VIs/Regularization-Methods-for-HVIs., and is run on Intel Xeon Platinum 8380 CPUs.
@@ 5.1 Equilibrium selection
This matches Problem (P) with $\mathsf{G}=\nabla\phi$ for the upper-level operator and $\mathsf{M}=\mathsf{F}+\operatorname{\mathsf{NC}}_{X}$ for the lower-level operator.
@@ Results.
Moreover, we note that the optimality and feasibility gaps may be computed through a small subproblem in this simple example. Figures 3 and 4 plot these values for the different initial points. Although we observe differences, we note that all curves are decreasing to 0. We note the importance of the tuning of the proximal parameter $\alpha$. In the present example, a smaller value yields faster convergence. As will be evidenced in later simulations, this is not always the case.
@@ 5.2 Least-Norm Least-Squares
We test the algorithm on four randomly generated instances with different dimensions given by $(P,Q)\in\{(70,100),(100,200),(100,500),(300,500)\}$. For each instance, the matrix $A$ is generated as a low-rank (and sparse) matrix via $A=U_{1}\cdot U_{2}$, where $U_{1}\in\mathbb{R}^{P\times R}$ and $U_{2}\in\mathbb{R}^{R\times Q}$ are chosen such that each component follows a standard normal law, with $R=50$ to induce a low-rank structure. Moreover, to ensure the eigenvalues are not too large such that the step-size is not too small, we clip each singular value to the interval $[0,10]$. A vector $s\in\mathbb{R}^{Q}$ is randomly generated with $20$ non-zero entries according to a uniform law on $[0,10]$, and the observation vector $b$ is constructed as $b=As+\nu$, where $\nu$ is a normal random vector with small entries, namely a standard normal vector scaled by a factor $0.1$. During the random generation, we ensure that $z=A^{\dagger}b\in X$, such that the analytical solution is known to be $z$, where $A^{\dagger}$ is the Moore-Penrose inverse.
@@ Results.
Figures 7 and 8 show the evolution of the error sequence $\|w_{n}-z\|$ and the lower-level function gap sequence $\tfrac{1}{2}\|Aw_{n}-b\|^{2}-\tfrac{1}{2}\|Az-b\|^{2}$ across iterations. As expected, we observe a decrease for each problem. We note that the initial decrease for the error sequence is consequential and stagnates for a number of iterations afterward. In fact, for $\alpha=0.1$, the prescribed $2000$ iterations are not enough to observe a second decrease phase, whereas it is sufficient for $\alpha\in\{1,10\}$.
@@ 5.3 Image Inpainting
For the inner loop, we set $\theta_{k}\equiv\theta=0.75$ and consider $\tau_{k}\equiv\tau$ to be constantly the largest value satisfying Equation (3.1) across every outer iteration. We set $\beta_{n}=(n+1)^{-\eta}$ for $\eta=0.55$ and $\varepsilon_{n}=\overline{\varepsilon}\cdot(n+1)^{-2}$ for $\overline{\varepsilon}=2$. We run a total of $5000$ iterations, with an initial point $w_{0}=Y_{\text{corrupt}}$. We set the regularization parameter to be $\sigma=50$, and generate $\Omega$ randomly such that $20\%$ of the image is corrupted.
@@ Results.
Figure 9 shows the original image, the corrupt image, and the image restored using the above procedure for various proximal parameters. Table 1 reports the total number of inner iterations run, along with the time spent to produce the results. Figure 10 plots the lower-level objective across outer iterations for each value of $\alpha$, which is expected to converge to the lower-level objective value and not to 0, as seen in the plot. We observe a loss in the solution quality for $\alpha=100$, although it is the fastest method. We also note that the solution qualities are very similar for $\alpha=0.1$ and $\alpha=1$, but the latter requires a larger number of total iterations.
@@
Figure 10: Evolution of lower-level objective function $\tfrac{1}{2}\|\mathcal{R}(w_{n})-Y_{\text{corrupt}}\|^{2}+\sigma\|w_{n}\|_{*}$ for various proximal parameters.
@@ 6 Conclusion
In this work, we propose a double-loop path-following method to solve hierarchical variational inequalities. The inner loop aims to solve an auxiliary problem and, given a fixed-point encoding of said problem, employs a Krasnoselskii-Mann-type iteration. This step is very flexible regarding the chosen encoding strategy, such that our method includes various known algorithms. The outer loop takes the form of a restarting procedure, giving rise to a diagonal equilibrium tracking method (DANTE). We obtain weak convergence of the averaged iterates as well as convergence rates, which are comparable to the state-of-the-art results for specific numerical schemes.
We believe that our work opens up possibilities for various interesting directions for future research. First, it would be of practical interest to devise a fully adaptive scheme without restarts. Second, an important extension is the inclusion of stochastic data. These directions are the subject of ongoing research.
Acknowledgements
Daniel Cortild acknowledges the support of the Clarendon Funds Scholarships. Meggie Marschner and Mathias Staudigl acknowledge financial support from the PGMO and the Deutsche Forschungsgemeinschaft (DFG) - Projektnummer 556222748 "non-stationary hierarchical minimization".